how to find the perimeter if the ratio is length 5 & width 3 and the area is 240 square
Answer by Theo(13342)

if the ratio of length to width is 5/3, then you get the following:

L/W = 5/3, where L stands for length and W stands for width.

solve for L to get L = 5W/3

The area of the rectangle is equal to 240.

formula for area is L*W = A

formula becomes L*W = 240

replace L with 5W/3 to get:

5W/3*W = 240

simplify to get 5W^2/3 = 240

multiply both sides of equation by 3 to get:

5W^2 = 240*3

divide both sides of equation by 5 to get:

W^2 = 240*3/5

simplify to get W^2 = 144

take the square root of both sides of equation to get W = 12

In equation of L = 5W/3, replace W with 12 to get:

L = 5*12/3

solve for L to get L = 20

you have L = 20 and W = 12

10*12 = 240 so the numbers are good because you know the area is 240.

you want to find the perimeter.

perimeter = 2L + 2W which becomes 2*20 + 2*12 which becomes 40 + 24 which becomes 64.

That's your solution.

Perimeter = 64 units.